# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This work has been submitted to IEEE Transactions on Information Forensics and Security

### v2

@@ I Introduction
As deep learning becomes dominant in many important areas, ensuring robustness becomes increasingly important. Deep neural networks are known to be vulnerable to adversarial attacks: small imperceptible perturbations in the inputs leading to incorrect decisions [1, 2]. Although many works have proposed heuristic defenses for training robust classifiers, they are often shown to be inadequate against adaptive attacks [3, 4, 5]. Therefore, a growing literature on certifiable robustness has emerged [6, 7, 8]; where the classifier’s prediction must be provably robust around any input within a perturbation budget.
@@
Another line of works [20, 21] propose using models for which each individual layer is $1$-Lipschitz. By enforcing orthogonal or near-orthogonal weight matrices, these networks naturally limit their sensitivity to input perturbations, contributing to a form of robustness that does not solely depend on Lipschitz constant estimation. This approach can complement the limitations of both Lipschitz constrained training and RS, particularly in handling larger models where direct Lipschitz estimation becomes impractical. Orthogonal constraints help maintain the global Lipschitz constant close to $1$.
@@
TABLE I: Comparison of certified test accuracy (%) on MNIST, CIFAR-10, and ImageNet under $\ell_{2}$ norm perturbation (see full results in Section IV). Each entry lists the certified accuracy using numbers taken from respective papers (RS results on MNIST follow from previous benchmark papers [22, 12]).
@@
Interestingly, this approach yields state-of-the-art results for several datasets. For instance, Table I compares the certified test accuracy of SPLITZ and existing state-of-art techniques for various values of $\epsilon$ (perturbation budget or certified radius) on the MNIST, CIFAR-10 and ImageNet datasets. For $\epsilon$ as large as $2.0$, where the state-of-the-art accuracy is $48.3\%$, SPLITZ achieves certified accuracy of around $62.3\%$. In Section IV, we present comprehensive set of results on MNIST, CIFAR-10 and ImageNet datasets. In addition, we also provide a detailed ablation study, and study the impact of various hyperparameters (such as the location of the split, randomized smoothing parameters, effects of local Lipschitz constant).
@@ III SPLITZ: Inference, Certification and Training
Figure 3: Comparison of certified radius with $\ell_{2}$ norm perturbation w.r.t RS [7], SmoothAdv [15], Smoothmix [12] and SPLITZ (ours), when varying levels of Gaussian noise $\sigma\in\{0.25,0.5,1.0\}$ on the CIFAR-10 dataset. We can observe that SPLITZ consistently outperforms RS [7] under different noise levels. For instance, when Radius = 1.0, SPLITZ achieves $43.2\%$ certified test accuracy while the best certified test accuracy of RS under different noise levels is $22.0\%$. We refer the reader to Table III for a comprehensive comparison of SPLITZ with several other approaches on CIFAR-10 dataset.
@@
where $\lambda\in[0,1]$ is a hyperparameter controlling the tradeoff between accuracy and robustness, $\theta$ is a learnable parameter to optimize the local Lipschitz constant, and $\text{Loss}(\cdot)$ is the loss function (e.g., cross entropy loss). Following the literature on randomized smoothing, we replace the expectation operators with their empirical estimates, and our loss function is:
@@
6:     Optimize $\gamma^{*}$ (discussed in Remark $1$) and calculate the corresponding local Lipschitz bound on $f^{*}_{L}$. (Eq. 18).
@@
where $W_{s}$ is the weight matrix of layer $s$. We include more details with respect to local Lipschitz bound in Appendix B.
@@ IV Experiments and Evaluation Results
MNIST Certified Test Accuracy at $\epsilon~{}(\%)$
Method Extra data 1.50 1.75 2.00 2.25 2.50
RS [7] 67.3 46.2 32.5 19.7 10.9
MACER [16] 73.0 50.0 36.0 28.0 -
Consistency [22] 82.2 70.5 45.5 37.2 28.0
SmoothMix [12] 81.8 70.7 44.9 37.1 29.3
DRT [23] 83.3 69.6 48.3 40.3 34.8
RS+OrthoNN [21] 70.1 49.7 33.2 21.0 10.5
SPLITZ 80.2 71.3 62.3 51.7 40.5
TABLE II: Comparison of certified test accuracy (%) on MNIST under $\ell_{2}$ norm perturbation. Each entry lists the certified accuracy using numbers taken from respective papers (RS results follow from previous benchmark papers [22, 12]).
@@
TABLE III: Comparison of the approximate certified test accuracy (%) on CIFAR-10 under $\ell_{2}$ norm perturbation. Extra data indicates whether their models incorporate other datasets in their models. Each entry lists the certified accuracy using numbers taken from respective papers.
Dataset Configuration For the MNIST and CIFAR-10 dataset(s), we draw $n_{0}=10^{2},~{}n_{1}=10^{5}$ noise samples to certify the smoothing model following [7, 14, 12]. For ImageNet dataset, we draw $n_{0}=10^{2},~{}n_{1}=10^{4}$ noise samples to certify the smoothing model following [7, 14, 12]. We set the Lipschitz threshold (see Sec III) as $\theta=0.5$. For local Lipschitz constrained training, we set tradeoff parameter $\lambda$ (see Sec III) evenly decrease from $0.8-0.4$, $0.7-0.5$ and $0.9-0.7$ respectively. We use one Nvidia P100 GPU to train the SPLITZ model with batch size $512,256,128$ respectively. We apply Adam Optimizer for three datasets. For the MNIST dataset, we conduct training for 150 epochs and utilize an early-stop strategy to search for the optimal classifier over an additional 150 epochs. We set the initial learning rate as 0.001. The learning rate is decayed (multiplied by 0.1) by 0.1 at every 50 epochs (50th, 100th…). For the CIFAR-10 dataset, we train 200 epochs using the ResNet110 and use the early-stop strategy to search for the optimal classifier over an additional 200 epochs. Furthermore, we set the initial learning rate as 0.001 and final learning rate as $10^{-6}$. The learning rate starts to evenly decay at each epoch from the half of the training epochs. For the ImageNet dataset, we train 200 epochs for the ResNet50 and set the initial learning rate as 0.01. The learning rate starts to decay at each 40 epochs. Our code is available at: https://anonymous.4open.science/r/SPLITZ-3C66. We report our more experimental details in Appendix C.
Baseline mechanisms We compare our method with various existing techniques proposed for robust training of smoothed classifiers, as listed below: (a) PixelDP [8]: cerified robust training with differential privacy mechanism; (b) RS [7]: standard randomized smoothing with the classifier trained with Gaussian augmentation; (c) SmoothAdv [15]: adversarial training combined with randomized smoothing; (d) MACER [16]: a regularization approach which maximizes the approximate certified radius; (e) Consistency [22]: a KL-divergence based regularization that minimizes the variance of smoothed classifiers $f(x+\delta)$ across $\delta$; (f) SmoothMix [12]: training on convex combinations of samples and corresponding adversarial on smoothed classifier; (g) Boosting [17]: a soft-ensemble scheme on smooth training; (h) DRT [23]: a lightweight regularized training on robust ensemble ML models; (i) ACES [39]: a selection-mechanism combined with a smoothed classifier; (j) DDS [14]: a denoised diffusion mechanism combined with a smoothed classifier; (k) RS [7] + OrthoNN [21]: an orthogonal convolutional layer followed by a randomized smoothing model.
@@ IV-A Main Results
Results on MNIST As showed in Table II, we can observe that SPLITZ outperforms other state-of-art approaches in almost every value of $\epsilon$. Impressively, we find that the SPLITZ classifier has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=2.50$, SPLITZ classifier achieves 40.5% compared to state-of-art top-1 test accuracy 34.8% certified test accuracy on the MNIST dataset. Moreover, when we increase $\epsilon$ from 1.50 to 2.50, RS drops from 67.3% to 10.9% decreasing 56.4% test accuracy. SPLITZ, however, maintains higher certified test accuracy from 80.2% to 40.5% maintaining relatively higher test accuracy.
Results on CIFAR10 As shown in Table III and Fig 3, our method outperforms the state-of-art approaches for every value of $\epsilon$ on CIFAR-10 dataset. Interestingly, we find that the SPLITZ training has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=1.0$, the model achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%. One hypothesis is that minimizing the Lipscitz bound of $f_{L}$ ($L_{f_{L}}\leq 1$) is able to boost the certified radius of the model. Intuitively, more samples are correctly classified while corresponding radius are larger than given $\epsilon$. In addition, we can observe the similar trend as MNIST dataset. SPLITZ maintains higher certified test accuracy when we increase $\epsilon$ from 0.25 to 1.00 compared to other state-of-art mechanisms.
@@
SPLITZ (ours) 43.2 38.6 31.2 23.9
TABLE IV: Comparison of the approximate certified test accuracy (%) on ImageNet under $\ell_{2}$ norm perturbation. The columns and rows have the same meaning as in Table III.
Results on ImageNet We show the comparison of different certified robustness techniques on ImageNet dataset in Table IV. We observe similar trends to MNIST and CIFAR10 datasets, where SPLITZ is effective on certified robustness with a wide range of image datasets. For instance, when $\epsilon=2.0$, SPLITZ achieve $31.2\%$ while the state-of-the-art have $30.4\%$. Moreover, SPLITZ consistently achieves better ACR (average certified radius) than other mechanisms in the ImageNet dataset, as shown in Table V.
Results on Average Certified Radius (ACR) We investigate the performance of SPLITZ using average certified radius (ACR), where we measure the correct samples’ average certified radius over the test datasets (MNIST, CIFAR-10, ImageNet). As shown in Table V, we provide the comprehensive comparison results of average certified radius (ACR) compared to other certified robust techniques. Our SPLITZ consistently outperforms others. For instance, when $\sigma=0.50$, the ACR of SPLITZ is 2.059 respectively, where the state-of-the-art is 0.933 in MNIST dataset. In addition, we can observe the same trends in CIFAR-10 and ImageNet dataset, where SPLITZ consistently outperforms the state-of-the-art when $\sigma=0.50,1.00$.
Empirical Robust Test Accuracy We study the empirical robust test accuracy in Table VI under $\ell_{2}$ FGSM, PGD and Auto-attack using RS and SPLITZ model with variance $\sigma=1$. We can observe that the SPLITZ consistently outperforms the RS method. For example, when $\epsilon=2.0$ under auto-attack, SPLITZ achieve the empirical test accuracy $23.15\%$ while RS is $7.72\%$.
IV-B Training and certifying time
Our SPLITZ model needs to vary the value of $\gamma$ (the size of the ball around input $x$) during the training epoch. Thus, we may need relatively more time to obtain the optimal model. To solve this, we apply the early stop mechanism to obtain a better optimized model during the training process. At the same time, we decay our learning rates during the training process. We use one Nvidia P100 GPU to train our SPLITZ model and report our training time (certifying time) for each one epoch (one image) in the Table VII.
@@ IV-B Training and certifying time
TABLE V: Comparison of average certified radius (ACR) across three different datasets (MNIST, CIFAR-10 and ImageNet). We can observe that for two datasets, SPLITZ consistently achieves better results compared to other state-of-art mechanisms. * is reported by [22, 12]
@@
TABLE VI: Comparison of empirical certified test accuracy under attacks with respect to CIFAR-10 dataset.
@@ IV-C Ablation Study
Training
Datasets Architecture Training per epoch (s) Certifying per image (s)
MNIST LeNet 15.1 1.0
CIFAR-10 ResNet110 59.5 50.7
ImageNet ResNet50 4810.5 83.1
TABLE VII: SPLITZ training time and certifying time across three datasets (MNIST using LeNet, CIFAR-10 using ResNet 110 and ImageNet using ResNet50).
@@
TABLE VIII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying the splitting layer on MNIST dataset with LeNet.
@@
Impact of splitting location As mentioned in Section III, our SPLITZ classifier can be optimized over different split ways, where we conduct the experiments and show our results in Table VIII. For example, when $\epsilon=2$, splitting after the $1^{st}$, or $2^{nd}$, or $3^{rd}$ layer result in certified accuracy of $62.3$%, $51.0$% and $49.4$% respectively. These results indicate that splitting the neural network early achieves better performance. Intuitively, splitting the neural network early helps the model minimize the local Lipschitz bound, which improves the certified robustness leading to a higher certified test accuracy given the same $\epsilon$. As the splitting becomes ”deeper”, estimating the local Lipshitz constant also becomes harder, which implies that a looser bound leads to smaller certified radius.
@@
Comparison of the orthogonal neural network (OrthoNN) and the SPLITZ In the baseline RS+OrthoNN model, during the training process, we enforce the first convolution layer of the classifier’s left half to act as an orthogonal convolution neural network. This is achieved by incorporating an orthogonal constraint regularization into the loss function. For the SPLITZ model, we impose a constraint on the local Lipschitz constant of the classifier’s left half. The comparison between RS+orthoNN and SPLITZ is illustrated in Table II with respect to the MNIST dataset, Table III with respect to the CIFAR-10 dataset and Table IV with respect to the ImageNet dataset. Our observations indicate that enforcing an orthogonal convolution layer on the classifier’s left half has a comparatively lesser impact than training with a local Lipschitz constraint.
@@
TABLE IX: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $Lip_{\theta}$ (the threshold of the local Lipschitz bound around input $x$) on MNIST dataset.
@@
TABLE X: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $\gamma$ (the size of the ball around input $x$) on MNIST dataset.
Effect of $\theta$ (Lipschitz threshold) As shown in Fig 4(b) and Table IX, we analyze the effect of the training threshold $\theta$ (See Eq III). For smaller values of $\epsilon$, SPLITZ with higher Lipschitz constant achieves better performances on accuracy. Conversely, SPLITZ with a smaller Lipschitz constant can boost certified radius, which obtains a relative higher certified test accuracy when $\epsilon$ is larger. These observations suggest a trade-off between the Lipschitz constant and certified accuracy. Thus, identifying the optimal Lipschitz threshold is essential for achieving a balances between robustness and utility. This ablation study further validates that the essence of our SPLITZ classifier lies in maintaining a relatively optimized Lipschitz constant for the left half of the classifier.
Effect of $\gamma$ (size of radius around input $x$) According to above results, constraining the local Lipschitz constant achieves better performance. To further explore the benefit of local Lipschitz constrained training, it is necessary to explore the indicator matrix $I^{v}$ in Eq 18, which depends on the size of the ball around the input (i.e., the hyperparameter $\gamma$). In Table X, we show how varying training $\gamma$ impacts the certified test accuracy for different values of $\epsilon$. We observe that smaller values of training $\gamma$ lead to higher certified accuracy for all values of $\epsilon$.

### v3

@@ I Introduction
As deep learning becomes dominant in many important areas, ensuring robustness becomes increasingly important. Deep neural networks are known to be vulnerable to adversarial attacks: small imperceptible perturbations in the inputs leading to incorrect decisions [1, 2]. Although many works have proposed heuristic defenses for training robust classifiers, they are often shown to be inadequate against adaptive attacks [3, 4, 5]. Therefore, a growing literature on certifiable robustness has emerged [6, 7, 8, 9, 10, 11]; where the classifier’s prediction must be provably robust around any input within a perturbation budget.
@@
Another line of works [23, 24] propose using models for which each individual layer is 111-Lipschitz. By enforcing orthogonal or near-orthogonal weight matrices, these networks naturally limit their sensitivity to input perturbations, contributing to a form of robustness that does not solely depend on Lipschitz constant estimation. This approach can complement the limitations of both Lipschitz constrained training and RS, particularly in handling larger models where direct Lipschitz estimation becomes impractical. Orthogonal constraints help maintain the global Lipschitz constant close to 111.
@@
TABLE I: Comparison of certified test accuracy (%) on MNIST, CIFAR-10, and ImageNet under $\ell_{2}$ norm perturbation ( see CIFAR-10 and ImageNet results in Section IV, and MNIST results in Appendix C). Each entry lists the certified accuracy using numbers taken from respective papers (RS results on MNIST follow from previous benchmark papers [25, 15]).
@@
[…]
@@ III SPLITZ: Inference, Certification and Training
[…]
@@
[…]
@@
6:  Optimize $\gamma^{*}$ (discussed in Remark 111) and calculate the corresponding local Lipschitz bound on $f^{*}_{L}$. (Eq. 18).
@@
where $W_{s}$ is the weight matrix of layer $s$. The local Lipschitz constant $L_{f_{L}}^{(\gamma)}(x)$ is computed as a product of spectral norms of weight matrices, modulated by activation-dependent indicator functions, which capture the network’s sensitivity to input perturbations. The bias parameters do not influence the Jacobian, as it is computed with respect to the input $x$, and thus have no effect on Equation (18). We include more details with respect to local Lipschitz bound in Appendix B.
@@ IV Experiments and Evaluation Results
Dataset Configuration For the MNIST and CIFAR-10 dataset(s), we draw $n_{0}=10^{2},~n_{1}=10^{5}$ noise samples to certify the smoothing model following [7, 17, 15]. For ImageNet dataset, we draw $n_{0}=10^{2},~n_{1}=10^{4}$ noise samples to certify the smoothing model following [7, 17, 15]. We set the Lipschitz threshold (see Sec III) as $\theta=0.5$. For local Lipschitz constrained training, we set tradeoff parameter $\lambda$ (see Sec III) evenly decrease from 0.80.40.8-0.40.8 - 0.4, 0.70.50.7-0.50.7 - 0.5 and 0.90.70.9-0.70.9 - 0.7 respectively. We use one Nvidia P100 GPU to train the SPLITZ model with batch size 512,256,128512,256,128512 , 256 , 128 respectively. We apply Adam Optimizer for three datasets. For the MNIST dataset, we conduct training for 150 epochs and utilize an early-stop strategy to search for the optimal classifier over an additional 150 epochs. We set the initial learning rate as 0.001. The learning rate is decayed (multiplied by 0.1) by 0.1 at every 50 epochs (50th, 100th…). For the CIFAR-10 dataset, we train 200 epochs using the ResNet110 and use the early-stop strategy to search for the optimal classifier over an additional 200 epochs. Furthermore, we set the initial learning rate as 0.001 and final learning rate as $10^{-6}$. The learning rate starts to evenly decay at each epoch from the half of the training epochs. For the ImageNet dataset, we train 200 epochs for the ResNet50 and set the initial learning rate as 0.01. The learning rate starts to decay at each 40 epochs. We use Adam optimizer for all datasets. We report our more experimental details in Appendix C.
Baseline mechanisms We compare our method with various existing techniques proposed for robust training of smoothed classifiers, as listed below: (a) PixelDP [9]: cerified robust training with differential privacy mechanism; (b) RS [7]: standard randomized smoothing with the classifier trained with Gaussian augmentation; (c) SmoothAdv [18]: adversarial training combined with randomized smoothing; (d) MACER [19]: a regularization approach which maximizes the approximate certified radius; (e) Consistency [25]: a KL-divergence based regularization that minimizes the variance of smoothed classifiers $f(x+\delta)$ across $\delta$; (f) SmoothMix [15]: training on convex combinations of samples and corresponding adversarial on smoothed classifier; (g) Boosting [20]: a soft-ensemble scheme on smooth training; (h) DRT [26]: a lightweight regularized training on robust ensemble ML models; (i) ACES [42]: a selection-mechanism combined with a smoothed classifier; (j) DDS [17]: a denoised diffusion mechanism combined with a smoothed classifier; (k) RS [7] + OrthoNN [24]: an orthogonal convolutional layer followed by a randomized smoothing model; (l) APNDC [43]: a novel diffusion classifier that acts as an ensemble of exact Posterior noised diffusion classifiers, while incurring no additional computational overhead.
@@
TABLE II: Comparison of the approximate certified test accuracy (%) on CIFAR-10 under $\ell_{2}$ norm perturbation. Extra data indicates whether their models incorporate other datasets in their models. Each entry lists the certified accuracy using numbers taken from respective papers.
@@ IV-A Main Results
Results on CIFAR10 As shown in Table II and Fig 3, our method outperforms the state-of-art approaches for every value of $\epsilon$ on CIFAR-10 dataset. Interestingly, we find that the SPLITZ training has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=1.0$, the model achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%. One hypothesis is that minimizing the Lipscitz bound of $f_{L}$ ($L_{f_{L}}\leq 1$) is able to boost the certified radius of the model. Intuitively, more samples are correctly classified while corresponding radius are larger than given $\epsilon$. In addition, we can observe the similar trend as MNIST dataset. SPLITZ maintains higher certified test accuracy when we increase $\epsilon$ from 0.25 to 1.00 compared to other state-of-art mechanisms.
@@
SPLITZ 43.2 38.6 31.2 23.9
TABLE III: Comparison of the approximate certified test accuracy (%) on ImageNet under $\ell_{2}$ norm perturbation. The columns and rows have the same meaning as in Table II.
@@
TABLE IV: Comparison of average certified radius (ACR) across three different datasets (MNIST, CIFAR-10 and ImageNet). We can observe that for two datasets, SPLITZ consistently achieves better results compared to other state-of-art mechanisms. * is reported by [25, 15]
[…]
Results on Average Certified Radius (ACR) We investigate the performance of SPLITZ using average certified radius (ACR), where we measure the correct samples’ average certified radius over the test datasets (MNIST, CIFAR-10, ImageNet). As shown in Table IV, we provide the comprehensive comparison results of average certified radius (ACR) compared to other certified robust techniques. Our SPLITZ consistently outperforms others. For instance, when $\sigma=0.50$, the ACR of SPLITZ is 2.059 respectively, where the state-of-the-art is 0.933 in MNIST dataset. In addition, we can observe the same trends in CIFAR-10 and ImageNet dataset, where SPLITZ consistently outperforms the state-of-the-art when $\sigma=0.50,1.00$.
@@
TABLE V: Comparison of empirical certified test accuracy under attacks with respect to CIFAR-10 dataset.
[…]
IV-B Training and certifying time
Our SPLITZ model needs to vary the value of $\gamma$ (the size of the ball around input $x$) during the training epoch. Thus, we may need relatively more time to obtain the optimal model. To solve this, we apply the early stop mechanism to obtain a better optimized model during the training process. At the same time, we decay our learning rates during the training process. We report the training time of SPLITZ and other baselines using one NVIDIA P100 GPU for CIFAR-10 and four NVIDIA P100 GPUs for ImageNet in the Table VI. We observe that SPLITZ requires slightly more training time than RS, but significantly less time than SmoothAdv and MACER.
@@ IV-C Ablation Study
Datasets Method Training per epoch (s)
CIFAR-10 RS 31.4
SmoothAdv 1990.1
MACER 504.0
SPLITZ 59.5
ImageNet RS 2154.5
SmoothAdv 7723.8
MACER 3537.1
SPLITZ 3160.5
TABLE VI: Comparison of SPLITZ training time across two datasets (CIFAR-10 using ResNet 110 and ImageNet using ResNet50) with RS [7], SmoothAdv [18] and MACER [19].
@@
TABLE VII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying the splitting layer on MNIST dataset with LeNet.
@@
Impact of splitting location As mentioned in Section III, our SPLITZ classifier can be optimized over different split ways, where we conduct the experiments and show our results in Table VII. For example, when $\epsilon=2$, splitting after the $1^{st}$, or $2^{nd}$, or $3^{rd}$ layer result in certified accuracy of 62.362.362.3%, 51.051.051.0% and 49.449.449.4% respectively. These results indicate that splitting the neural network early achieves better performance. Intuitively, splitting the neural network early helps the model minimize the local Lipschitz bound, which improves the certified robustness leading to a higher certified test accuracy given the same $\epsilon$. As the splitting becomes ”deeper”, estimating the local Lipshitz constant also becomes harder, which implies that a looser bound leads to smaller certified radius.
@@
Comparison of the orthogonal neural network (OrthoNN) and the SPLITZ In the baseline RS+OrthoNN model, during the training process, we enforce the first convolution layer of the classifier’s left half to act as an orthogonal convolution neural network. This is achieved by incorporating an orthogonal constraint regularization into the loss function. For the SPLITZ model, we impose a constraint on the local Lipschitz constant of the classifier’s left half. The main advantage of SPLITZ is that it splits the neural network into two halves and constrains the Lipschitz constant of the left part to be less than 1 (instead of exactly 1 while using OrthoNN), which can boost the certified radius and enables a better trade-off between robustness and accuracy. The comparison between RS+orthoNN and SPLITZ is illustrated in Table X with respect to the MNIST dataset, Table II with respect to the CIFAR-10 dataset and Table III with respect to the ImageNet dataset. Our observations indicate that enforcing an orthogonal convolution layer on the classifier’s left half has a comparatively lesser impact than training with a local Lipschitz constraint.
@@
TABLE VIII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $Lip_{\theta}$ (the threshold of the local Lipschitz bound around input $x$) on MNIST dataset.
@@
TABLE IX: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $\gamma$ (the size of the ball around input $x$) on MNIST dataset.
@@
Effect of $\theta$ (Lipschitz threshold) As shown in Fig 4(b) and Table VIII, we analyze the effect of the training threshold $\theta$ (See Eq III). For smaller values of $\epsilon$, SPLITZ with higher Lipschitz constant achieves better performances on accuracy. Conversely, SPLITZ with a smaller Lipschitz constant can boost certified radius, which obtains a relative higher certified test accuracy when $\epsilon$ is larger. These observations suggest a trade-off between the Lipschitz constant and certified accuracy. Thus, identifying the optimal Lipschitz threshold is essential for achieving a balances between robustness and utility. This ablation study further validates that the essence of our SPLITZ classifier lies in maintaining a relatively optimized Lipschitz constant for the left half of the classifier.
Effect of $\gamma$ (size of radius around input $x$) According to above results, constraining the local Lipschitz constant achieves better performance. To further explore the benefit of local Lipschitz constrained training, it is necessary to explore the indicator matrix $I^{v}$ in Eq 18, which depends on the size of the ball around the input (i.e., the hyperparameter $\gamma$). In Table IX, we show how varying training $\gamma$ impacts the certified test accuracy for different values of $\epsilon$. We observe that smaller values of training $\gamma$ lead to higher certified accuracy for all values of $\epsilon$.